# Bayesian Modular Inference for Copula Models with Potentially Misspecified Marginals

### 12 Mar 2026 v1

@@ I.1 Proofs
$p_{\text{SMI},\gamma}(\psi,u,\eta,\widetilde{\eta}\mid\mathcal{D})=p_{\text{SMI},\gamma}(\psi,u,\widetilde{\eta}\mid\mathcal{D})p(\eta\mid\psi,\mathcal{D}),$
which yields
Assumption 1.
For some $\varepsilon>0$, $\gamma_{j}\in[\varepsilon,1-\varepsilon]$ for all $j=1,\dots,d$.
@@
Define the $u$-marginal SMI posterior
@@
$=\prod_{i=1}^{n}\Delta^{\gamma}_{i}(\widetilde{\eta},\psi)$
$=\Delta^{\gamma}_{n}(\widetilde{\eta},\psi).$
The allows us to write the posterior $p_{\text{SMI},\gamma}(\psi\mid\mathcal{D})$ in the more convenient form: for $f_{n}(\widetilde{\eta}):=\prod_{i=1}^{n}\prod_{j=1}^{d}f_{j}(y_{ij}\mid\widetilde{\eta}_{j})$, we have
$p_{\text{SMI},\gamma}(\psi,\widetilde{\eta}\mid\mathcal{D})=\frac{\Delta^{\gamma}_{n}(\widetilde{\eta},\psi)p(\psi,{\eta})f_{n}(\widetilde{\eta})}{\int\int\Delta^{\gamma}_{n}(\eta,\psi)p(\psi,\widetilde{\eta})f_{n}(\widetilde{\eta})\mathrm{d}\eta\mathrm{d}\psi},$
which also defines the $\psi$-marginal SMI posterior
$p_{\text{SMI},\gamma}(\psi\mid\mathcal{D})=\frac{\int\Delta^{\gamma}_{n}(\eta,\psi)p(\psi,{\eta})f_{n}({\eta})\mathrm{d}\eta}{\int\int\Delta^{\gamma}_{n}(\eta,\psi)p(\psi,{\eta})f_{n}({\eta})\mathrm{d}\eta\mathrm{d}\psi}.$
@@
$\mathsf{L}^{\gamma}_{n}(\varphi)=-[\log\Delta^{\gamma}_{n}(\widetilde{\eta},\psi)+\log f_{n}(\widetilde{\eta})]/n,$
we have that
$p_{\text{SMI},\gamma}(\varphi\mid\mathcal{D})=\frac{\exp\{-n\mathsf{L}^{\gamma}_{n}(\varphi)\}p(\varphi)}{\int\exp\{-n\mathsf{L}^{\gamma}_{n}(\varphi)\}p(\varphi)\mathrm{d}\varphi}\equiv\frac{\exp\{-n[\mathsf{L}^{\gamma}_{n}(\varphi)-\mathsf{L}^{\gamma}_{n}(\varphi^{\gamma}_{0})]\}p(\varphi)}{\int\exp\{-n[\mathsf{L}^{\gamma}_{n}(\varphi)-\mathsf{L}^{\gamma}_{n}(\varphi^{\gamma}_{0})]\}p(\varphi)\mathrm{d}\varphi}.$ (I.1)
@@
(i) For any $\varphi\in\Phi$, the function $\varphi\mapsto\mathsf{L}(\varphi;\gamma):=\mathbb{E}\mathsf{L}_{n}^{\gamma}(\varphi)$ exists. (ii) For some $\varepsilon>0$, and any $\gamma\in[0,1-\varepsilon]^{d}$, there exist $\gamma\mapsto\varphi^{\gamma}_{0}$ such that, for any $\delta>0$,
$\inf_{\varphi:\|\varphi-\varphi_{0}(\gamma)\|>\delta}\mathsf{L}(\varphi;\gamma)\geq\mathsf{L}(\varphi^{\gamma}_{0};\gamma).$
(iii) For some $C>0$ and $\alpha>0$, $\|\varphi-\varphi_{0}^{\gamma}\|\leq C[\mathsf{L}(\varphi;\gamma)-\mathsf{L}(\varphi^{\gamma}_{0};\gamma)]^{\alpha}$.
@@
There exists a positive sequence $r_{n}\downarrow 0$ as $n\rightarrow\infty$ such that, for any $\gamma\in[\varepsilon,1-\varepsilon]^{d}$,
$\mathbb{E}\left[\sup_{\varphi\in\Phi}|\mathsf{L}_{n}^{\gamma}(\varphi)-\mathbb{E}[\mathsf{L}_{n}^{\gamma}(\varphi)]|\right]\leq r_{n}$
@@
Let $\epsilon>0$ with $n\epsilon\geq 1$, and let $\alpha>0$ be as in Assumption 2. For $\mathcal{A}_{\epsilon}:=\{\theta\in\Theta:|\mathsf{L}(\varphi;\gamma)-\mathsf{L}(\varphi^{\gamma}_{0};\gamma)|\leq\epsilon^{\alpha}\}$, the prior satisfies $\int_{\Phi}\mathds{1}\{\varphi\in\mathcal{A}_{\epsilon}\}p(\varphi)\mathrm{d}\varphi\geq e^{-n\epsilon^{\alpha}}$.
@@
Under Assumption 1-4, for $n$ large enough,
$\mathbb{E}\int\mathds{1}\left\{\|\psi-\psi_{0}^{\gamma}\|>Mr_{n}\right\}p_{\mathrm{SMI},\gamma}(\psi\mid\mathcal{D})\mathrm{d}\psi\leq 1/M$
@@
Let us simplify notation and drop dependence on $\gamma$ for now. Similarly, let us write $q_{n}(\varphi)=p_{\mathrm{SMI},\gamma}(\varphi\mid\mathcal{D})$, and $q_{n}(\psi)=\int q_{n}(\varphi)\mathrm{d}\eta$. The proof proceeds through two main steps.
@@
    First, we derive a bound on the probability of the joint set $\|\varphi-\varphi_{0}\|>M\epsilon$ under $q_{n}(\varphi)$.
@@
    We then use a marginalization argument to arrive at a similar bound for $q_{n}(\psi)$.
Step 1. Fix $\epsilon>0$. Apply Assumption 2 (iii) and Markov’s inequality to obtain
$Q_{n}\left[\|\varphi-\varphi_{0}\|>M\epsilon\right]$ $=Q_{n}\left[\|\varphi-\varphi_{0}\|^{1/\alpha}>(M\epsilon)^{1/\alpha}\right]$
$\leq Q_{n}\left[\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\}>\frac{(M\epsilon)^{1/\alpha}}{C^{1/\alpha}}\right]$
$\lesssim\frac{1}{(M\epsilon)^{1/\alpha}}\int\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\}q_{n}(\varphi)\mathrm{d}\varphi$ (I.2)
Rewrite and bound the integral term in (I.2) as
@@
$\leq 2\sup_{\varphi\in\Phi}|\mathsf{L}(\varphi)-\mathsf{L}_{n}(\varphi)|+\int\left\{\mathsf{L}_{n}(\varphi)-\mathsf{L}_{n}(\varphi_{0})\right\}q_{n}(\varphi)\mathrm{d}\varphi$ (I.3)
Substituting (I.3) into (I.2) and taking expectations we obtain
$\mathbb{E}Q_{n}\left[\|\varphi-\varphi_{0}\|>M\epsilon\right]$ $\lesssim\frac{2}{(M\epsilon)^{\frac{1}{\alpha}}}\mathbb{E}\left[\sup_{\varphi\in\Phi}|\left\{\mathsf{L}(\varphi)-\mathsf{L}_{n}(\varphi)\right\}|\right]+\frac{1}{(M\epsilon)^{\frac{1}{\alpha}}}\mathbb{E}\int\left\{\mathsf{L}_{n}(\varphi)-\mathsf{L}_{n}(\varphi_{0})\right\}q_{n}(\varphi)\mathrm{d}\varphi$
$\lesssim\frac{r_{n}}{(M\epsilon)^{\frac{1}{\alpha}}}+\frac{1}{(M\epsilon)^{\frac{1}{\alpha}}}\mathbb{E}\int\left\{\mathsf{L}_{n}(\varphi)-\mathsf{L}_{n}(\varphi_{0})\right\}q_{n}(\varphi)\mathrm{d}\varphi.$ (I.4)
Step 2. We now use the variational definition of $q_{n}(\varphi)$ to bound the second term in (I.4). Since $q_{n}(\varphi):=p_{\text{SMI},\gamma}(\varphi\mid\mathcal{D})$ as in (I.1), $q_{n}(\varphi)$ can be equivalently defined as the solution to the following variational optimization problem:
$q_{n}(\varphi):=\operatornamewithlimits{argmin}_{q\in\mathcal{P}(\Phi)}\left\{\int[\mathsf{L}_{n}(\varphi)-\mathsf{L}_{n}(\varphi_{0})]q(\varphi)\mathrm{d}\varphi+\frac{\mathrm{KL}(q\|p)}{n}\right\}.$ (I.5)
In particular, we see that, from (I.5),
@@
where the first inequality uses the fact that $\mathrm{KL}(q_{n}\|p)/\lambda>0$, and the second follows direction from the definition in (I.5). Apply expectations to both sides of the above, and use Jensen’s inequality to obtain:
@@
$\leq\inf_{q\in\mathcal{P}(\Phi)}\left\{\int\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\}q(\varphi)\mathrm{d}\varphi+\frac{\mathrm{KL}(q\|p)}{n}\right\}.$ (I.6)
@@
Since the above is true for all $q\in\mathcal{P}(\Phi)$, take $q=\rho_{n}$ in (I.6) and apply Assumption 4 to see that
@@
$\leq\epsilon^{\alpha}+\frac{\mathrm{KL}(\rho_{n}\|p)}{n},$
@@
$\mathrm{KL}(\rho_{n}\|p)=-\log\int_{\mathcal{A}_{\epsilon}}p(\varphi)\mathrm{d}\varphi\leq n\epsilon^{\alpha}.$
@@
$\mathbb{E}\int\{\mathsf{L}_{n}(\varphi)-\mathsf{L}_{n}(\varphi_{0})\}q_{n}(\varphi)\mathrm{d}\varphi\leq 2\epsilon^{\alpha}.$ (I.7)
Applying the bound in (I.7) into (I.4) we arrive at
$\mathbb{E}Q_{n}\left[\|\varphi-\varphi_{0}\|>M\epsilon\right]$ $\lesssim\frac{1}{M}\frac{r_{n}}{\epsilon^{\frac{1}{\alpha}}}+\frac{2}{M}$
Taking $\epsilon^{1/\alpha}r_{n}\geq 1$ implies that $\mathbb{E}Q_{n}\left[\|\varphi-\varphi_{0}\|>M\epsilon\right]\lesssim 1/M$.
Step 3. We can now transfer the bound obtained in Step 2 using the definition of $q_{n}(\psi)$. In particular, fix $\delta>0$ and take $\epsilon$ as in the proofs of Steps 1 and 2. For this choice, if $\|\varphi-\varphi_{0}\|\leq M\epsilon$, then $\|\psi-\psi_{0}\|\leq M\epsilon$ since $\psi$ is a subset $\varphi$. Therefore, we see that
$\{\psi:\|\psi-\psi_{0}\|>M\epsilon\}\subseteq\{\varphi:\|\varphi-\varphi\|>M\epsilon\}.$
@@
$\mathbb{E}Q_{n}[\|\psi-\psi_{0}\|>M\epsilon]$ $:=\mathbb{E}\int\int\mathds{1}[\|\psi-\psi_{0}\|>M\epsilon]q_{n}(\psi,\eta)\mathrm{d}\eta$
$\leq\mathbb{E}Q_{n}[\|\varphi-\varphi_{0}\|>M\epsilon]$
@@
from Steps 1 and 2. Taking $\delta=r_{n}$ yields the stated result. ∎
I.2 Primitive conditions for Assumption 3
Assumption 3 is a high-level condition that is somewhat hard to interpret. Herein, we show that Assumption 3 is satisfied by lower-level conditions that can be checked on a per-model basis. These conditions are as follows.

### 02 Apr 2026 v2

@@ B.1 Proofs
$p_{\text{SMI},\gamma}(\psi,u,\eta,\widetilde{\eta}\mid\mathcal{D})=p_{\text{SMI},\gamma}(\psi,u,\widetilde{\eta}\mid\mathcal{D})p(\eta\mid\psi,\mathcal{D}).$
@@
Define the SMI posterior, marginal of $u$,
@@
$=:\Delta_{n}(\widetilde{\eta},\psi;\gamma).$
This allows us to write the posterior $p_{\text{SMI},\gamma}(\psi\mid\mathcal{D})$ in a more convenient form: for $f_{n}(\widetilde{\eta}):=\prod_{i=1}^{n}\prod_{j=1}^{d}f_{j}(y_{ij}\mid\widetilde{\eta}_{j})$, we have
$p_{\text{SMI},\gamma}(\psi,\widetilde{\eta}\mid\mathcal{D})=\frac{\Delta_{n}(\widetilde{\eta},\psi;\gamma)p(\psi,{\eta})f_{n}(\widetilde{\eta})}{\int\int\Delta_{n}(\eta,\psi;\gamma)p(\psi,\widetilde{\eta})f_{n}(\widetilde{\eta})\mathrm{d}\widetilde{\eta}\mathrm{d}\psi},$
which defines the $\psi$-marginal SMI posterior as
$p_{\text{SMI},\gamma}(\psi\mid\mathcal{D})=\frac{\int\Delta_{n}(\widetilde{\eta},\psi;\gamma)p(\psi,\widetilde{\eta})f_{n}(\widetilde{\eta})\mathrm{d}\widetilde{\eta}}{\int\int\Delta_{n}(\widetilde{\eta},\psi;\gamma)p(\psi,{\eta})f_{n}(\widetilde{\eta})\mathrm{d}\widetilde{\eta}\mathrm{d}\psi}.$
@@
$p_{\text{SMI},\gamma}(\varphi\mid\mathcal{D})=\frac{\exp\{-n\mathsf{L}_{n}(\varphi;\gamma)\}p(\varphi)}{\int\exp\{-n\mathsf{L}_{n}(\varphi;\gamma)\}p(\varphi)\mathrm{d}\varphi}\equiv\frac{\exp\{-n[\mathsf{L}_{n}(\varphi;\gamma)-\mathsf{L}_{n}(\varphi_{0}(\gamma);\gamma)]\}p(\varphi)}{\int\exp\{-n[\mathsf{L}_{n}(\varphi;\gamma)-\mathsf{L}_{n}(\varphi_{0}(\gamma);\gamma)]\}p(\varphi)\mathrm{d}\varphi}.$ (B.1)
We maintain the following condition that stipulates the support of the influence parameter $\gamma$.
@@
For all $j=1,\dots,d$, $\gamma_{j}\in[0,1]$, with $\gamma_{j}\not\in\{0,1\}$ for at least one $j$.
In addition to the requirement on $\gamma$ maintained in Assumption 1, we require the following regularity conditions.
@@
(i) For any $\varphi\in\Phi:=\Psi\times\mathcal{E}$, the function $\varphi\mapsto\mathsf{L}(\varphi;\gamma):=\mathbb{E}\mathsf{L}_{n}(\varphi;\gamma)$ exists. (ii) For some $\varepsilon>0$, and any $\gamma$ as in Assumption 1, there exist $\gamma\mapsto\varphi_{0}(\gamma)$ such that, for all $\delta>0$, and some $\delta^{\prime}>0$,
$\inf_{\varphi:\|\varphi-\varphi_{0}^{\gamma}\|>\delta}\mathsf{L}(\varphi;\gamma)>\mathsf{L}(\varphi_{0}(\gamma);\gamma)+\delta^{\prime}.$
(iii) There exist $C>0$, $\alpha>0$, and $U\subset\Phi$ containing $\varphi_{0}(\gamma)$ such that, for all $\varphi\in U$, $\|\varphi-\varphi_{0}^{\gamma}\|\leq C[\mathsf{L}(\varphi;\gamma)-\mathsf{L}(\varphi_{0}(\gamma);\gamma)]^{\alpha}$.
@@
There exists a positive sequence $r_{n}\downarrow 0$ as $n\rightarrow\infty$ such that, for any $\gamma$ as in Assumption 1,
$\mathbb{E}\left[\sup_{\varphi\in\Phi}|\mathsf{L}_{n}(\varphi;\gamma)-\mathbb{E}[\mathsf{L}_{n}(\varphi;\gamma)]|\right]\leq r_{n}.$
Assumption 4.
Let $\epsilon>0$ with $n\epsilon\geq 1$. For $\mathcal{A}_{\epsilon}:=\{\varphi\in\Phi:|\mathsf{L}(\varphi;\gamma)-\mathsf{L}(\varphi_{0}(\gamma);\gamma)|\leq\epsilon\}$, the prior satisfies $\int_{\Phi}\mathds{1}\{\varphi\in\mathcal{A}_{\epsilon}\}p(\varphi)\mathrm{d}\varphi\geq e^{-n\epsilon}$.
Assumption 2(i)-(ii) requires that a minimizer of the loss $\mathsf{L}(\varphi;\gamma)$ exists for any $\gamma$ at which we wish to compute the SMP; while Assumption 2(iii) is a type of Holder-continuity that we must maintain to link fluctuations in the loss near the minimizer, to differences in the parameter space. Assumption 3 is a high-level condition that requires the empirical process between the sample and population loss converges at some rate; sufficient conditions for this are discussed in Section B.2. Assumption 4 is a standard prior mass condition used in most applications of generalized Bayesian inference.
Theorem 1 in the main text is a direct consequence of Theorems 1 and 2 that follow.
@@
Under Assumption 1-4, for $n$ and $M>0$ large enough,
$\mathbb{E}\int\mathds{1}\left\{\|\psi-\psi_{0}(\gamma)\|>CM^{\alpha}r_{n}^{\alpha}\right\}p_{\mathrm{SMI},\gamma}(\psi\mid\mathcal{D})\mathrm{d}\psi\leq 1/M.$
@@
Let us simplify notation and drop dependence on $\gamma$ for now. Similarly, let us write $q_{n}(\varphi)=p_{\mathrm{SMI},\gamma}(\varphi\mid\mathcal{D})$, and $q_{n}(\psi)=\int q_{n}(\varphi)\mathrm{d}\widetilde{\eta}$. The proof proceeds through two main steps.
@@
    First, we derive a bound on the probability of $\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\}>M\epsilon$ under $q_{n}(\varphi)$.
@@
    We then transfer this bound to $\|\varphi-\varphi_{0}\|$.
  4. 4.
    A marginalization argument is then used to arrive at a similar bound for $q_{n}(\psi)$.
Step 1. Fix $\epsilon>0$. From Markov’s inequality,
$Q_{n}\left[\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\}>{(M\epsilon)}\right]\leq\frac{1}{(M\epsilon)}\int\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\}q_{n}(\varphi)\mathrm{d}\varphi.$ (B.2)
Rewrite and bound the integral term in (B.2) as
@@
$\leq 2\sup_{\varphi\in\Phi}|\mathsf{L}(\varphi)-\mathsf{L}_{n}(\varphi)|+\int\left\{\mathsf{L}_{n}(\varphi)-\mathsf{L}_{n}(\varphi_{0})\right\}q_{n}(\varphi)\mathrm{d}\varphi$ (B.3)
Substituting (B.3) into (B.2), taking expectations and using Assumption 3 we obtain
$\mathbb{E}Q_{n}\left[\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\}>M\epsilon\right]$
$\leq\frac{2}{(M\epsilon)}\mathbb{E}\left[\sup_{\varphi\in\Phi}|\left\{\mathsf{L}(\varphi)-\mathsf{L}_{n}(\varphi)\right\}|\right]+\frac{1}{(M\epsilon)}\mathbb{E}\int\left\{\mathsf{L}_{n}(\varphi)-\mathsf{L}_{n}(\varphi_{0})\right\}q_{n}(\varphi)\mathrm{d}\varphi$
$\leq\frac{r_{n}}{(M\epsilon)}+\frac{1}{(M\epsilon)}\mathbb{E}\int\left\{\mathsf{L}_{n}(\varphi)-\mathsf{L}_{n}(\varphi_{0})\right\}q_{n}(\varphi)\mathrm{d}\varphi.$ (B.4)
Step 2. We now use the variational definition of $q_{n}(\varphi)$ to bound the second term in (B.4). Since $q_{n}(\varphi):=p_{\text{SMI},\gamma}(\varphi\mid\mathcal{D})$ as in (B.1), $q_{n}(\varphi)$ can be equivalently defined as the solution to the following variational optimization problem:
$q_{n}(\varphi):=\operatornamewithlimits{argmin}_{q\in\mathcal{P}(\Phi)}\left\{\int[\mathsf{L}_{n}(\varphi)-\mathsf{L}_{n}(\varphi_{0})]q(\varphi)\mathrm{d}\varphi+\frac{\mathrm{KL}(q\|p)}{n}\right\}.$ (B.5)
In particular, we see that, from (B.5),
@@
where the first inequality uses the fact that $\mathrm{KL}(q_{n}\|p)>0$, and the equality follows (B.5). Apply expectations to both sides of the above, and use Jensen’s inequality to obtain:
@@
$\leq\inf_{q\in\mathcal{P}(\Phi)}\left\{\int\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\}q(\varphi)\mathrm{d}\varphi+\frac{\mathrm{KL}(q\|p)}{n}\right\}.$ (B.6)
@@
Since the above is true for all $q\in\mathcal{P}(\Phi)$, take $q=\rho_{n}$ in (B.6) and apply Assumption 4 to see that
@@
$\leq\epsilon+\frac{\mathrm{KL}(\rho_{n}\|p)}{n},$
@@
$\mathrm{KL}(\rho_{n}\|p)=-\log\int_{\mathcal{A}_{\epsilon}}p(\varphi)\mathrm{d}\varphi\leq n\epsilon.$
@@
$\mathbb{E}\int\{\mathsf{L}_{n}(\varphi)-\mathsf{L}_{n}(\varphi_{0})\}q_{n}(\varphi)\mathrm{d}\varphi\leq 2\epsilon.$ (B.7)
Applying the bound in (B.7) into (B.4) we arrive at
$\mathbb{E}Q_{n}\left[\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\}>M\epsilon\right]$ $\leq\frac{1}{M}\frac{r_{n}}{\epsilon}+\frac{2}{M}=\frac{r_{n}+2\epsilon}{M\epsilon}$
Take $\epsilon=r_{n}$ so that $\frac{r_{n}+2\epsilon}{M\epsilon}\leq 4/M$ implies that $\mathbb{E}Q_{n}\left[\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\}>Mr_{n}\right]\leq 4/M\asymp 1/M$.
Step 3. Let $\delta>0$ be such that $\{\varphi\in\Phi:\|\varphi-\varphi_{0}\|\leq\delta\}\subset U$, where $U$ is defined in Assumption 2(iii). By Assumption 2(ii), there exists $\delta^{\prime}>0$ such that $\|\varphi-\varphi_{0}\|>\delta$ implies that $\mathsf{L}(\varphi)>\mathsf{L}(\varphi_{0})+\delta^{\prime}$. Let $n$ be large enough so that $\delta^{\prime}>Mr_{n}$, which implies that
$\left\{\varphi\in\Phi:\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\}\leq Mr_{n}\right\}\subset U.$
However, from Step 3 we know that
$\mathbb{E}Q_{n}\left\{\mathsf{L}(\varphi)-\mathsf{L}(\varphi_{0})\leq Mr_{n}\right\}\gtrsim 1-\frac{1}{M},$
and applying Assumption 2(iii) implies that
$\mathbb{E}Q_{n}\left\{\|\varphi-\varphi_{0}\|\leq CM^{\alpha}r_{n}^{\alpha}\right\}\gtrsim 1-\frac{1}{M}.$
Negating the event implies, for $M$ large enough, $\mathbb{E}Q_{n}\left\{\|\varphi-\varphi_{0}\|>CM^{\alpha}r_{n}^{\alpha}\right\}\lesssim\frac{1}{M}.$
Step 4. We can now transfer the bound obtained in Step 3 using the definition of $q_{n}(\psi)$. In particular, fix $\delta>0$ and $\delta^{\prime}$ as in Step 3. For this choice, if $\|\varphi-\varphi_{0}\|\leq CM^{\alpha}r^{\alpha}_{n}$, then $\|\psi-\psi_{0}\|\leq CM^{\alpha}r^{\alpha}_{n}$ since $\psi$ is a subset $\varphi$. Therefore, we see that
$\{\psi:\|\psi-\psi_{0}\|>CM^{\alpha}r^{\alpha}_{n}\}\subseteq\{\varphi:\|\varphi-\varphi\|>CM^{\alpha}r^{\alpha}_{n}\}.$
@@
$\mathbb{E}Q_{n}[\|\psi-\psi_{0}\|>CM^{\alpha}r^{\alpha}_{n}]$ $:=\mathbb{E}\int\int\mathds{1}[\|\psi-\psi_{0}\|>CM^{\alpha}r^{\alpha}_{n}]q_{n}(\psi,\eta)\mathrm{d}\eta$
$\leq\mathbb{E}Q_{n}[\|\varphi-\varphi_{0}\|>CM^{\alpha}r^{\alpha}_{n}]$
@@
where the last line follows by applying the conclusion of Step 3. ∎
Theorem 2.
Under Assumptions 1-4,
$\int\int\mathds{1}\{\|\eta-\eta_{0}^{\gamma}\|>CM^{\alpha}_{n}r_{n}^{\alpha}\}p_{\mathrm{SMI},\gamma}(\eta,\psi\mid\mathcal{D})\mathrm{d}\psi\mathrm{d}\eta$
converges to zero in probability.
Proof.
First, recall that $p_{\mathrm{SMI},\gamma}(\eta,\psi\mid\mathcal{D})=p_{\mathrm{SMI},\gamma}(\psi\mid\mathcal{D})p(\eta\mid\psi,\mathcal{D})$. We again drop terms dependence on $\gamma$ to simply notation. Fix $\varepsilon>0$, and separate the integral in question into:
$\int\int\mathds{1}\{\|\eta-\eta_{0}\|>\varepsilon\}p_{\mathrm{SMI},\gamma}(\eta,\psi\mid\mathcal{D})\mathrm{d}\psi\mathrm{d}\eta$
$=\int\int_{\|\psi-\psi_{0}\|>CM^{\alpha}_{n}r_{n}^{\alpha}}\mathds{1}\{\|\eta-\eta_{0}\|>\varepsilon\}p_{\mathrm{SMI},\gamma}(\eta,\psi\mid\mathcal{D})\mathrm{d}\psi\mathrm{d}\eta$
$+\int\int_{\|\psi-\psi_{0}\|\leq CM^{\alpha}_{n}r_{n}^{\alpha}}\mathds{1}\{\|\eta-\eta_{0}\|>\varepsilon\}p_{\mathrm{SMI},\gamma}(\eta,\psi\mid\mathcal{D})\mathrm{d}\psi\mathrm{d}\eta.$ (B.8)
The first term satisfies the following trivial bound: for $n$ and $M$ large enough,
$\int\int_{\|\psi-\psi_{0}\|>CM^{\alpha}_{n}r_{n}^{\alpha}}\mathds{1}\{\|\eta-\eta_{0}\|>\varepsilon\}p_{\mathrm{SMI},\gamma}(\eta,\psi\mid\mathcal{D})\mathrm{d}\psi\mathrm{d}\eta$
$\leq\int_{\|\psi-\psi_{0}\|>CM^{\alpha}_{n}r_{n}^{\alpha}}p_{\mathrm{SMI},\gamma}(\psi\mid\mathcal{D})\mathrm{d}\psi$
$\leq 1/M,$ (B.9)
with probability converging to one by Theorem 1.
For the second term, we note that Theorem 1 implies that $p_{\mathrm{SMI},\gamma}(\psi\mid\mathcal{D})\Rightarrow\delta_{\psi_{0}}$, and since $p(\eta\mid\psi,\mathcal{D})$ is bounded for all $\psi$, we have
$\int\int_{\|\psi-\psi_{0}\|\leq CM^{\alpha}_{n}r_{n}^{\alpha}}\mathds{1}\{\|\eta-\eta_{0}\|>\varepsilon\}p_{\mathrm{SMI},\gamma}(\eta,\psi\mid\mathcal{D})\mathrm{d}\psi\mathrm{d}\eta$
$\Rightarrow\int\mathds{1}\{\|\eta-\eta_{0}\|>\varepsilon\}p(\eta\mid\psi_{0},\mathcal{D})\mathrm{d}\eta.$ (B.10)
However,
$p(\eta\mid\psi_{0},\mathcal{D})\propto p(\eta\mid\psi_{0})p_{\mathrm{SMI},\gamma=1}(\mathcal{D}\mid\eta,\psi_{0})p(\eta\mid\varphi_{0})\propto e^{-n\mathsf{L}_{n}(\eta,\varphi_{0})}p(\eta\mid\varphi_{0})$
Since Assumptions 1-4 are satisfied for all $\varphi=(\psi^{\top},\eta^{\top})^{\top}$, they remain satisfied at $(\psi_{0}^{\top},\eta^{\top})^{\top}$.
Hence, repeating similar arguments to those in Theorem 1 demonstrates that, for any $\varepsilon>0$, and $n$ and $M$ large enough,
$\mathbb{E}\int\mathds{1}\{\|\eta-\eta_{0}\|>\varepsilon\}p(\eta\mid\psi_{0},\mathcal{D})\mathrm{d}\eta\lesssim 1/M.$
Taking $\varepsilon=CM^{\alpha}r_{n}\alpha$ we have that, with probability converging to one,
$\int\mathds{1}\{\|\eta-\eta_{0}\|>\varepsilon\}p(\eta\mid\psi_{\star},\mathcal{D})\mathrm{d}\eta\leq 1/M.$
Now, for $M=M_{n}\rightarrow\infty$ as $n\rightarrow\infty$, possibly slowly, we can apply the above into equation (B.1) to conclude that
$\int\int_{\|\psi-\psi_{0}\|\leq CM^{\alpha}_{n}r_{n}^{\alpha}}\mathds{1}\{\|\eta-\eta_{0}\|>\varepsilon\}p_{\mathrm{SMI},\gamma}(\eta,\psi\mid\mathcal{D})\mathrm{d}\psi\mathrm{d}\eta\leq 1/M_{n}$ (B.11)
with probability converging to one.
Applying (B.9) and (B.11) into equation (B.8) delivers the stated convergence. ∎
B.2 Primitive conditions for Assumption 3
Assumption 3 is a high-level condition that is somewhat hard to interpret. Herein, we show that Assumption 3 is satisfied by lower-level conditions on the marginal models and copula function which can be checked on a per-model basis.